OpenStudy (anonymous):

the surface area of a rectangular prism is found using the formula SA=2(lw+lh+wh) where l is length, w is width, and h is height. if the surface area of a rectangle prism is 324m2, with length 5 m and width 12 m what is the height in meters?

OpenStudy (anonymous):

\[ 324=2((5)(12)+(5)h+(12)h) \]

OpenStudy (anonymous):

Simplify and solve for \(h\).

OpenStudy (anonymous):

i am just not getting it

OpenStudy (anonymous):

I'll do it step by step then.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[ SA=2(lw+lh+wh) \] "the surface area of a rectangle prism is 324m2" \[ SA= 324\] \[ 324=2(lw+lh+wh) \]

OpenStudy (anonymous):

Does this step make sense?

OpenStudy (anonymous):

kind of

OpenStudy (anonymous):

Why doesn't it make sense?

OpenStudy (anonymous):

ok i get it thanks

OpenStudy (anonymous):

\[ 324=2(lw+lh+wh) \]"with length 5 m"\[ l=5 \] \[ 324=2((5)w+(5)h+wh) \]

OpenStudy (anonymous):

Do you get this step?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

width 12 \[w=12\] \[ 324=2((5)(12)+(5)h+(12)h) \]

OpenStudy (anonymous):

Okay, now are you ready to solve for \(h\)?